# Mapping High-Performance Regions in Battery Scheduling across Data Uncertainty, Battery Design, and Planning Horizons

### 12 Apr 2026 v1

@@ Abstract
This study presents a triadic analysis of energy storage operation under multi-stage model predictive control, investigating the interplay between data characteristics, forecast uncertainty, planning horizon, and battery c-rate. Synthetic datasets are generated to systematically explore variations in data profiles and uncertainty, enabling parametrization and the construction of relationships that map these characteristics to optimal horizon length. Results reveal the presence of an effective horizon, defined as the look-ahead length beyond which additional forecast information provides limited operational benefit. Accounting for this horizon can reduce computational costs while maintaining optimal performance. The study provides optimal horizon lengths across a broad range of combinations of battery types, uncertainty levels, and data profiles, offering practical guidance for industrial storage operation. It also quantifies revenue losses due to forecast uncertainty, showing that errors can impact performance even for fast batteries. Finally, the framework lays the groundwork for future machine learning approaches that map dataset parametrization to optimal horizons, supporting continuous optimization in industrial settings without heavy computation.
@@ 1 Introduction
This observation motivates the concept of an effective planning horizon: the portion of the nominal optimization horizon over which future information can meaningfully influence scheduling decisions, given the battery’s physical constraints. Importantly, this effective planning horizon emerges endogenously from the interaction between battery dynamics and the temporal structure of the input data, and persists even under perfect foresight. In this sense, fast-cycling batteries exhibit a self-limiting effect on the optimization window, naturally reducing exposure to long-term forecast uncertainty .
@@
This paper, provides a methodology to analyse the interplay between battery power–capacity ratios, data characteristics, and optimization horizon length (see Figure 1). The study is based on a deterministic multistage scheduling framework, to deliberately isolate the structural effects of horizon length without confounding influences from stochastic or risk-aware solution methods. Specifically, the methodology aims to (i) investigate how battery design influences the ability to exploit temporal variations in forecast data, (ii) examine how data properties (such as amplitude variability, seasonality patterns, and forecast error growth) affect the value of longer horizons, and (iii) identify high-performance regions in the combined space of battery characteristics, data profiles, and planning horizons. The resulting insights should provide guidance for selecting battery designs and optimization horizons adapted to the nature of the underlying data, and establish a foundation for future extensions incorporating stochastic optimization or other optimization under uncertainty approaches.
@@ 3.1.1 Synthetic ground truth data
Ground-truth time series are generated using a hybrid parametric–stochastic construction that combines a Fourier-inspired deterministic component with a stochastic time-series model. The deterministic backbone is defined as a superposition of a finite number of sinusoidal components:
@@
Alternative approaches for synthetic data generation include state-space models Durbin and Koopman (2001); Kalman (1960), or data-driven generative models including variational autoencoders Kingma and Welling (2014); Fortuin et al. (2020) and generative adversarial networks Goodfellow et al. (2014); Yoon et al. (2019). While these methods can produce realistic samples, they typically offer less transparent parameterizations and reduced interpretability. In contrast, the hybrid approach adopted here preserves the intuitive, low-dimensional structure of Fourier-based representations while augmenting them with stochastic dynamics, enabling a balance between interpretability and realism. This is particularly important for establishing mappings between data characteristics, battery parameters, and optimal planning horizons.
@@ 4.1.2 Synthetic forecasts
Forecast series are generated following the procedure described in the Methods section. For each synthetic dataset, rolling forecasts with a horizon of 72 hours are produced at intervals of 3 hours; that is, a new 72-hour hourly forecast is issued every 3 hours. All the datasets were also tested for a publishing interval of 6 hours. However, to avoid unnecessary duplication of results, only the 3-hour interval is presented in the results, while the 6-hour case is deferred to the Discussion section. For each of the nine synthetic time-series instances (three families with three variants each), five corresponding forecast series are generated. Every combination is then tested for 6 different batteries, yielding a total of 9 x 5 x 6 = 270 experiments, each of them tested for 45 different planning horizons (8,910 2-week hourly resolution optimization runs).
With regards to the decay of forecast accuracy (increase in the standard deviation of the error distribution), a simple linear approach was chosen to maintain consistency across datasets and ensure interpretability. Based on empirical observations from the real datasets, an exponential increase was deemed unrealistic, as error growth exhibited patterns closer to linear or weakly seasonal behavior. In particular, the day-ahead market showed some intra-day structure, with higher variability during peak hours and lower variability overnight, whereas the mFRR data did not exhibit a clear seasonal pattern. The linear choice provides a simple and transparent representation of forecast degradation while avoiding additional assumptions about intra-day structure.
@@
As mentioned in the Methods section, the variation between these forecasts is introduced through the width of the error distribution. Specifically, the standard deviation of the forecast error is multiplied by ‘uncertainty factors’ (u.f.) of 0.1, 1, 3, 6, and 10, producing five forecast series with progressively increasing levels of uncertainty. These values have been adjusted (for interpretability) in such a way that the uncertainty factor $u.f.=1.0$ corresponds, roughly, to the uncertainty obtained from real forecasts in the day-ahead market for a span of 24h. While the uncertainty factor $u.f.=10.0$ is comparable to the level uncertainty found in the mFRR market for the same period, as illustrated in Fig. 8.
@@ 6.3 Revenue degradation and uncertainty gap
The uncertainty gap appears in most scenarios, and it tends to become more pronounced as forecast uncertainty increases, as expected. Although the relative losses observed in the experiments can be inferred visually from Figures 131415, and 16 for experiments with rolling forecasts published every 3 hours, and from Figures 242526, and 27 in the Appendix for 6-hour publication intervals, some representative numerical examples of revenue loss are also provided below for reference.
@@
    Second, the reported losses within the uncertainty gap correspond to reductions in top-line revenue. In realistic economic settings, where operating expenditure (OPEX) and capital expenditure (CAPEX) must be recovered from these revenues, even moderate percentage losses can have a large impact on net profitability. This especially true when profit margins are tight. For illustration, one may consider a simple case where a battery generates 100 EUR of daily revenue under optimal conditions. A 5% reduction corresponds to 95 EUR of revenue. Assuming fixed daily costs of 80 EUR, profit would decrease from 20 EUR to 15 EUR, representing a 25% reduction in profit despite only a 5% drop in revenue.
@@ 7 Conclusions and future work
This work has introduced a methodological framework for analysing planning horizon selection in rolling-horizon optimization under forecast uncertainty, with the aim of providing interpretable and practically relevant insights for energy system operation. A central design choice of the framework is the explicit parameterization of the information structure, as well as other aspects like forecast publication intervals or the rolling-horizon strides. By treating these elements as exogenous and configurable parameters rather than fixed assumptions, the approach remains applicable across a wide range of operational settings and lays the groundwork for future models capable of inferring optimal horizons directly from time-series parameters.
@@
  • Existence of an effective planning horizon. The empirical results consistently indicate the presence of an effective planning horizon, defined as the point beyond which additional forecast information does not lead to further changes in revenue. In practical terms, this corresponds to a plateau in the revenue curve, where extending the horizon yields negligible gains. This observation suggests that the effective horizon is a structural property of the interaction between the signal and the battery system, and it motivates future work aimed at formally establishing its existence and characterising it mathematically.
@@ 7.1 Future work
  • Improved modelling of forecast uncertainty: The synthetic forecasts used in this study allow for large deviations under high uncertainty levels, which may influence some observed effects (e.g., extended effective horizons at extreme uncertainty). Future work should consider more realistic uncertainty models, including calibrated probabilistic forecasts and empirically derived error distributions.

### 06 May 2026 v2

@@ Abstract
This study presents a controlled parametric framework for analysing energy storage planning under uncertainty in a multi-stage model predictive control setting. The framework enables a broad and systematic exploration through parametrized generation of synthetic datasets in the context of energy price arbitrage. It facilitates the study of the joint effects of battery characteristics, signal structure, forecast uncertainty, and planning horizon on revenue performance in energy storage optimisation, which are rarely considered together. The analysis is driven by two objectives. First, it characterises how these interacting factors influence operational revenue and its sensitivity to planning horizon selection, including economic losses caused by deviations from optimal horizons. This provides guidance on expected horizon ranges and their impact on revenue and computational cost. Second, it enables a compact parametrization of the relationships between battery properties, data characteristics, forecast uncertainty, and horizon-dependent performance, providing a basis for future modelling of optimal planning horizon length.
Results show that the framework captures consistent structural dependencies across configurations and provides meaningful guidance for horizon selection under uncertainty. In particular, increasing forecast uncertainty systematically reduces the optimal planning horizon across battery types, reflecting the diminishing value of long-term information under increasingly unreliable forecasts. Comparison with real market data shows that the parametrization reproduces the main qualitative trends of optimal horizon behaviour, suggesting its potential as a lightweight surrogate for more complex simulation-based analysis. Overall, the study provides an interpretable framework for understanding the interaction between battery properties, signal structure, forecast uncertainty, and planning horizon in energy storage optimisation.
  • A controlled parametric framework is developed to study energy storage operation under forecast uncertainty in a multistage MPC setting.
  • Synthetic datasets enable systematic analysis of how battery characteristics, data structure, forecast uncertainty and horizon length jointly affect revenue performance.
  • A consistent reduction of the optimal planning horizon length is observed with increasing forecast uncertainty across battery types.
  • A compact parametrization of battery–data–uncertainty interactions provides a basis for future modelling of optimal horizon length.
  • Comparison with real market data shows that the framework captures key qualitative trends in horizon-dependent performance.
Keywords: energy storage optimisation; battery scheduling optimization; model predictive control; optimal planning horizon; forecast uncertainty; battery arbitrage; synthetic data; parametric modelling
@@ 1 Introduction
This observation reveals a saturation behaviour in the marginal value of extending the optimization horizon, whereby additional forecast information eventually ceases to meaningfully affect scheduling decisions. This motivates the concept of an effective planning horizon: the optimization horizon beyond which future information no longer influences scheduling decisions, given the battery’s physical constraints. Importantly, this effective planning horizon emerges endogenously from the interaction between battery dynamics and the temporal structure of the input data, and persists even under perfect foresight. In this sense, fast-cycling batteries exhibit a self-limiting effect on the optimization window, naturally reducing exposure to long-term forecast uncertainty .
@@
This paper, provides a methodology to analyse the interplay between battery power–capacity ratios, data characteristics, forecast uncertainty and optimization horizon length (see Figure 1). The study is based on a deterministic multistage scheduling framework, to deliberately isolate the structural effects of horizon length without confounding influences from stochastic or risk-aware solution methods. Specifically, the methodology aims to (i) investigate how battery design influences the ability to exploit temporal variations in forecast data, (ii) examine how data properties (such as amplitude variability, seasonality patterns, and forecast error growth) affect the value of longer horizons, and (iii) identify high-performance regions in the combined space of battery characteristics, data profiles, and planning horizons. The resulting insights should provide guidance for selecting battery designs and optimization horizons adapted to the nature of the underlying data, and establish a foundation for future extensions incorporating stochastic optimization or other optimization under uncertainty approaches.
The remainder of this paper is structured as follows. Section 2 presents the literature review. Section 3 introduces the proposed methodology. Section 4 describes the experimental setup, including the generation of synthetic datasets and forecast scenarios. Section 5 presents the main results of the study. Section 6 discusses the findings and provides a comparison with real market data. Section 7 concludes the paper and outlines directions for future work. Additional results are provided in the Appendix.
@@ 3.1.1 Synthetic ground truth data
Ground-truth time series are generated using a hybrid parametric–stochastic construction that combines a Fourier-inspired deterministic component with a stochastic time series model. The deterministic backbone is defined as a superposition of a finite number of sinusoidal components:
@@
Alternative approaches for synthetic data generation include state-space models Durbin and Koopman (2001); Kalman (1960), or data-driven generative models including variational autoencoders Kingma and Welling (2014); Fortuin et al. (2020) and generative adversarial networks Goodfellow et al. (2014); Yoon et al. (2019). While these methods can produce realistic samples, they typically offer less transparent parametrizations and reduced interpretability. In contrast, the hybrid approach adopted here preserves the intuitive, low-dimensional structure of Fourier-based representations while augmenting them with stochastic dynamics, enabling a balance between interpretability and realism. This is particularly important for establishing mappings between data characteristics, battery parameters, and optimal planning horizons.
@@ 4.1.2 Synthetic forecasts
Forecast series are generated following the procedure described in the Methods section. For each synthetic dataset, rolling forecasts with a horizon of 72 hours are produced at intervals of 3 hours, i.e. a new 72-hour hourly forecast is issued every 3 hours. All datasets were also tested with a 6-hour publishing interval; however, only the 3-hour case is reported in the Results section, while the 6-hour case is deferred to the Appendix. For each of the nine synthetic time series instances (three families with three variants each), five corresponding forecast series are generated. Every combination is then evaluated across 6 different batteries, yielding $9\times 5\times 6=270$ experiments, each tested over 45 planning horizons (8,910 two-week hourly-resolution optimization runs).
The evolution of forecast uncertainty over the horizon is modelled as a linear increase in the standard deviation of the error distribution. This choice is motivated by the fact that the two markets considered in this study exhibit different temporal behaviours in forecast accuracy degradation. The linear formulation provides a simple and low-dimensional parametrisation that captures both cases within a reasonable level of fidelity, while preserving interpretability. In particular, the day-ahead market shows intra-day structure, with higher variability during peak hours and lower variability overnight, whereas the mFRR data does not exhibit a clear seasonal pattern. This linear representation avoids introducing additional assumptions about intra-day structure that are not consistently supported across both datasets.
@@
Forecast variability is controlled through uncertainty factors (u.f.) of 0.1, 1.0, 3.0, 6.0, and 10.0, which scale the standard deviation of the forecast error distribution, as described in the Methods section. These values are chosen such that $u.f.=1.0$ approximately matches the uncertainty level observed in real day-ahead forecasts over a 24-hour time span, while $u.f.=10.0$ corresponds approximately to the higher uncertainty regime observed in the mFRR market over the same horizon (see Fig. 8).
@@
This setup allows systematic evaluation of the impact of forecast uncertainty on revenue and optimal planning horizon length. Additional degrees of freedom within the error model, such as varying autocorrelation structure, are kept fixed in order to isolate the effect of variance scaling. Figure 9 illustrates an example forecast trajectory corresponding to $u.f.=1.0$ for the Fourier + 1.0 SARIMA instance of the day-ahead family. For clarity, only a single forecast realization is shown rather than the full rolling sequence.
@@ 6.3 Revenue degradation and uncertainty gap
The uncertainty gap appears in most scenarios, and it tends to become more pronounced as forecast uncertainty increases, as expected. Although the relative losses observed in the experiments can be inferred visually from Figures 131415, and 16 for experiments with rolling forecasts published every 3 hours, and from Figures 262728, and 29 in the Appendix for 6-hour publication intervals, some representative numerical examples of revenue loss are also provided below for reference.
@@
    Second, the reported losses within the uncertainty gap correspond to reductions in top-line revenue. In realistic economic settings, where operating expenditure (OPEX) and capital expenditure (CAPEX) must be recovered from these revenues, even moderate percentage losses can have a large impact on net profitability. This especially true when profit margins are tight. For illustration, one may consider a simple case where a battery generates 100 EUR/MWh of daily revenue under optimal conditions. A 5% reduction corresponds to 95 EUR/MWh of revenue. Assuming fixed daily costs of 80 EUR/MWh, profit would decrease from 20 to 15 EUR/MWh, representing a 25% reduction in profit despite only a 5% drop in revenue.
@@
6.3.1 Comparison with real data
The objective of this comparison is to evaluate whether the behavioural patterns and quantitative relationships observed in the synthetic experiments carry over to real-world data. The comparison assesses whether the parametrization scheme developed in this study can serve as a proxy for real data in approximating operational behaviour, and in particular, optimal horizon length and revenue loss due to uncertainty.
To enable this comparison, a representative subset of results consistent with the synthetic experimental setup is considered. Revenue-horizon curves are presented for battery cycle times of 1h, 2h, 4h, and 6h across both the real day-ahead and mFRR datasets, from which the synthetic ground truth signals were originally derived (Figures 24 and 25).
A quantitative summary of optimal planning horizons for both markets is provided in Table 8.
Battery cycle DA real / synthetic (h) mFRR real / synthetic (h)
1h 6 / 6 4 / 4
2h 7 / 9 4 / 6* (8)
4h 10 / 14* (19) 5 / 6
6h 18 / 9 5 / 6
* ‘Business’ optimal horizon: denotes a shorter planning horizon that produces revenue very close to the maximum, within a negligible difference.
Table 8: Comparison of optimal planning horizons between real data and synthetic experiments (Fourier + 1.0 SARIMA). Synthetic results correspond to $u.f.=1.0$ for DA and $u.f.=10.0$ for mFRR.
Day-ahead dataset
The day-ahead dataset exhibits broadly similar behaviour to the synthetic experiments, particularly for fast batteries. For 1h cycle batteries, the optimal horizon matches exactly between real and synthetic cases (6h vs. 6h), while for 2h batteries the difference remains moderate (7h vs. 9h). For the 4h battery, a larger discrepancy is observed between the real optimal horizon (10h) and the synthetic estimate (19h, with a ‘business’ optimum at 14h). However, the revenue curve in this region is relatively flat, and neighbouring horizons produce similar performance, indicating that this difference is less significant in operational terms than the raw values suggest. For the 6h battery, a more pronounced divergence is observed, with optimal horizons shifting from 18h in real data to approximately 9h in the synthetic setting. This difference can be attributed to the structure of the synthetic forecast errors. As described in the Methods section, the uncertainty factor was calibrated to match aggregate forecast error over a 24h horizon. However, the synthetic model imposes a simplified temporal structure with smooth degradation, whereas real day-ahead forecasts exhibit a more structured evolution of accuracy (e.g. daily seasonality effects). As a result, long-horizon accuracy is effectively underestimated in the synthetic setting, which reduces the value of extended planning horizons for slower batteries.
A similar pattern is observed in revenue degradation beyond the optimal horizon. Importantly, the presence of an uncertainty gap is consistently confirmed in the real data, with even stronger effects in some cases (e.g. up to 9.0% revenue drop for the 2h battery vs. 0.5%). However, for slower batteries, the degradation is less pronounced than in the synthetic experiments (Table 9), again consistent with better preservation of long-horizon forecast accuracy in real day-ahead data. Generally, the day-ahead results indicate that the synthetic framework captures the main qualitative trends, and provides a reasonable approximation of optimal horizon structure, particularly for fast batteries. Remaining discrepancies are primarily associated with the quantities in revenue loss caused by the temporal structure chosen to model forecast uncertainty.
mFRR dataset
In the mFRR case, the agreement between real and synthetic results is generally acceptable, particularly with respect to optimal horizon magnitude across battery types. As shown in Table 8, optimal horizons in real data are tightly clustered between 4h and 5h across all battery cycle times, while synthetic results produce slightly wider but still very comparable values. This suggests that horizon selection is driven primarily by short-term uncertainty structure rather than long-range error dynamics in this market, which consistent with its unpredictable nature.
The most notable difference, however, consists of a shift in the mapping between battery cycle times and the magnitude of revenue degradation beyond the optimal horizon. In the real data, the uncertainty gap leads to consistently strong losses, reaching up to 78.7% already for 2h batteries, and 43.7%, 34.6% for 4h and 6h cycles, respectively. While the synthetic experiments reproduce the same qualitative behaviour, similar levels of degradation are only observed for slower batteries, typically in the 4h–6h range, as shown in Table 9. This indicates that the underlying sensitivity to forecast errors is not fundamentally different between the two settings, but rather manifests at different effective time scales. In operational terms, real mFRR data appears to exhibit a compressed response in which faster batteries are exposed to effects that, in the synthetic parametrization, become similarly intense only for slower storage configurations. To explain the reason for this shifted behaviour, it is important to acknowledge a modeling issue in the synthetic forecasts: although the synthetic parametrization was calibrated to match aggregate error levels (e.g. MAE over 24h), it does not fully reproduce the rapid early degradation of forecast accuracy observed in real mFRR data, followed by a persistently low-accuracy regime.
Overall, the synthetic framework provides a reasonable approximation of real-world behaviour in both markets. Agreement is particularly strong in the day-ahead case for fast batteries, while the mFRR market shows good qualitative consistency in optimal horizon structure, despite differences in the temporal evolution of forecast uncertainty. These results suggest that the proposed parametrization captures first-order operational effects, while more detailed modelling of forecast dynamics would be required to fully reproduce market-specific behaviours.
Finally, in terms of revenue, all datasets exhibit a better performance aligned with the c-rate axis (faster batteries earn more revenue). Although, in the real datasets, this effect is less pronounced, especially in the case of the day-ahead dataset. This again may be attributed to the slower loss in accuracy over time of real day-ahead forecasts.
Battery cycle (h) DA loss (%) real / synthetic mFRR loss (%) real / synthetic
1h 0.2 / 0.2 6.4 / 0.1
2h 9.0 / 0.5 78.7 / 0.0
4h 5.4 / 0.2 43.7 / 26.1
6h 1.0 / 13.4 34.6 / 42.3
Table 9: Comparison of revenue losses beyond the optimal planning horizon between real data and synthetic experiments (Fourier + 1.0 SARIMA).
Figure 24: Day-ahead dataset with real ground truth and forecasts: revenue as a function of planning horizon length for four battery configurations with cycle times of 1h, 2h, 4h, and 6h (as indicated in the plots). All results are based on rolling forecasts with a 3-hour publishing interval.
Figure 25: mFRR dataset with real ground truth and forecasts: revenue as a function of planning horizon length for four battery configurations with cycle times of 1h, 2h, 4h, and 6h (as indicated in the plots). All results are based on rolling forecasts with a 3-hour publishing interval.
@@ 7 Conclusions and future work
This work has introduced a methodological framework for analysing planning horizon selection in rolling-horizon optimization under forecast uncertainty, with the aim of providing interpretable and practically relevant insights for energy system operation. A central design choice of the framework is the explicit parametrization of the information structure, as well as other aspects like forecast publication intervals or the rolling-horizon strides. By treating these elements as exogenous and configurable parameters rather than fixed assumptions, the approach remains applicable across a wide range of operational settings and lays the groundwork for future models capable of inferring optimal horizons directly from time series parameters.
@@
  • Existence of an effective planning horizon: The empirical results consistently indicate the presence of an effective planning horizon, defined as the point beyond which additional forecast information does not lead to further changes in revenue. In practical terms, this corresponds to a plateau in the revenue curve, where extending the horizon yields negligible gains. This observation suggests that the effective horizon is a structural property of the interaction between the signal and the battery system, and it motivates future work aimed at formally establishing its existence and characterising it mathematically.
@@
  • Comparison with real data: The comparison with real market data shows that the synthetic framework captures the main structural dependencies governing optimal planning horizons and revenue sensitivity to uncertainty. The results support the validity of the proposed parametrization scheme as a useful approximation tool for studying horizon selection, particularly in terms of identifying relative trends across battery types and uncertainty regimes.
    From a quantitative perspective, however, the comparison highlights the need for improved modelling of forecast uncertainty, in particular its temporal structure and accuracy decay over the time axis. Addressing this limitation is necessary for the framework to provide not only qualitative but also quantitatively reliable results, and remains an important direction for future work.
@@ 7.1 Future work
  • Improved modelling and calibration of forecast uncertainty: The synthetic forecasts used in this study do not fully capture the temporal structure of forecast accuracy observed in real data. In particular, the comparison with real markets indicates that uncertainty exhibits market-specific patterns that have been left out of the parametrization scheme. Future work should focus on improving the scheme so that captures these effects, while preserving a compact and interpretable set of parameters.